OpenStudy (nikkindavis):

What is the solution of the system? Use substitution. y = 3x + 10 y = 4x

OpenStudy (anonymous):

srry i would help but idk how to do tht

OpenStudy (nikkindavis):

Np

OpenStudy (snowsurf):

You substitute one of the equations into the other.

OpenStudy (nikkindavis):

I know I got an answer I just don't know if I am right

OpenStudy (snowsurf):

What is your answer?

OpenStudy (snowsurf):

Show your steps then I point out where you may went wrong

OpenStudy (nikkindavis):

I got (–10, –40)

OpenStudy (snowsurf):

the 10 is right but should be positive value. y = 3x + 10 y= 4x Since y is isolated for both equations you substitute 4x = 3x + 10 Solve for x 4x - 3x = 3x + 10 -3x x = 10